OpenStudy (kabase1234):

PLEASE HELP I'll fan and medal Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ...

OpenStudy (kabase1234):

@phi

OpenStudy (kabase1234):

@mayankdevnani

OpenStudy (mayankdevnani):

\[\large \bf a_n=a+(n-1)d\]

OpenStudy (mayankdevnani):

can you identify the value of a,n and d ?

OpenStudy (mayankdevnani):

where a=first term n=nth term d=common difference

OpenStudy (kabase1234):

Is -15 the first term?

OpenStudy (kabase1234):

And can you show me how to find the common difference again? I keep forgetting

OpenStudy (mayankdevnani):

yea! a=-15 remember :- common difference is difference between any two consecutive terms so, d=-6-(-15)=-6+15=9

OpenStudy (mayankdevnani):

understood?

OpenStudy (kabase1234):

Yes that makes so much sense thank you

OpenStudy (kabase1234):

So then the equation would look like a-15=a(-15-1)9?

OpenStudy (kabase1234):

@mayankdevnani

OpenStudy (kabase1234):

I got -7

OpenStudy (kabase1234):

Oh wait I didn't do the other side yet

OpenStudy (mayankdevnani):

\[\large \bf a_n=-15+(n-1)9\] simplify it

OpenStudy (kabase1234):

9n-24?

OpenStudy (kabase1234):

@yanasidlinskiy

OpenStudy (kabase1234):

Can you help me please

OpenStudy (kabase1234):

Choices an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1)

OpenStudy (phi):

maya posted the answer \[ a_n=-15+(n-1)9 \] the last term is (n-1) times 9 as you know, you can switch the order when multiplying example: 2*3 = 3*2 all your answers show the 9 first, but that is not problem: switch the order: \[ a_n=-15+9(n-1)\]